# Matrix Rosenthal and Concentration Inequalities for Markov Chains with Applications in Statistical Learning

### v1

@@ 2 Main Results
In this section, we present our main results: the Rosenthal-Burkholder inequality for matrix martingales, as well as matrix Rosenthal, Hoeffding, and Bernstein inequalities for ergodic Markov chains. We will also give dimension-free versions of these results.
We find that the dimension-free results can be applied to linear operator-valued martingales in infinite-dimensional separable Hilbert spaces. For brevity, we will not change the notation, and our dimension-free results still hold if ${\mathbb{H}}^{d}$ is replaced by the space of all self-adjoint operators on the separable Hilbert space $\mathcal{X}$}, and ${\mathbb{C}}^{d_{1}{\times}d_{2}}$ is replaced by the space of all bounded linear operators from the separable Hilbert space ${\mathcal{X}}_{1}$ to the separable Hilbert space ${\mathcal{X}}_{2}$.
@@ 2.1 The Rosenthal-Burkholder Inequality for Matrix Martingales
We denote by $({\bm{X}}_{k})_{k\in{\mathbb{N}}}$ the martingale difference sequence, that is, ${\bm{X}}_{k}={\bm{M}}_{k}-{\bm{M}}_{k-1}\in{\mathbb{H}}^{d}$ (${\bm{M}}_{-1}:=\bm{0}_{d{\times}d}$). And we denote by $(\left\langle{{\bm{M}}}\right\rangle_{k})_{k\in{\mathbb{N}}}$ the quadratic variation sequence, that is, $\left\langle{{\bm{M}}}\right\rangle_{k}=\sum_{j=0}^{k}{\mathbb{E}}_{j-1}{\bm{X}}_{j}^{2}\in{\mathbb{H}}^{d}_{+}$. Throughout this section, we assume that $\left\langle{{\bm{M}}}\right\rangle_{\infty}:=\lim_{k\to\infty}\left\langle{{\bm{M}}}\right\rangle_{k}<\infty$ element-wisely almost surely. As a result, by the martingale convergence theorem (Durrett, 2019, Theorem 4.5.2), $\lim_{k\to\infty}{{\bm{M}}}_{k}={\bm{M}}_{\infty}\in{\mathbb{H}}^{d}$ almost surely.
@@
Dimension-Free: As mentioned in Minsker (2017), in the case of ultra-high dimensions or even infinite dimensions, it is desirable for the results not to depend on the dimension $d$ of the ambient space. To this end, we resort to the notion of effective rank (Koltchinskii and Lounici, 2016). For any ${\bm{\Upsilon}}\in{\mathbb{H}}^{d}_{+}$, its effective rank is defined as
$r({\bm{\Upsilon}})=\frac{\operatorname{tr}({\bm{\Upsilon}})}{\left\|{{\bm{\Upsilon}}}\right\|}.$
One can check that $r({\bm{\Upsilon}})\leq d$ and $r({\bm{\Upsilon}})$ can be much smaller than $d$ if the eigenvalues of $\bm{\Upsilon}$} rapidly decay to 0.
@@
$\left\|{\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{k}\|}\right\|_{p}\leq C_{R,1}\sqrt{p\vee\log d}\Big{\|}\left\|{\left\langle{{\bm{M}}}\right\rangle_{\infty}}\right\|^{1/2}\Big{\|}_{p}+C_{R,2}(p\vee\log d)\left\|{\sup_{k\in{\mathbb{N}}}\|{\bm{X}}_{k}\|}\right\|_{p},$ (6)
where $C_{R,1}=87$, $C_{R,2}=50$. When $p\geq 117$, the constants can be smaller: $C_{R,1}=64$, $C_{R,2}=28$.
@@
    In the rectangular case (a), if we further assume there exist ${\bm{\Upsilon}}_{1}\in{\mathbb{H}}_{+}^{d_{1}}$, ${\bm{\Upsilon}}_{2}\in{\mathbb{H}}_{+}^{d_{2}}$ such that $\sum_{k=0}^{\infty}{\mathbb{E}}_{k-1}[{\bm{D}}_{k}{\bm{D}}_{k}^{*}]\preccurlyeq{\bm{\Upsilon}}_{1}$ and $\sum_{k=0}^{\infty}{\mathbb{E}}_{k-1}[{\bm{D}}_{k}^{*}{\bm{D}}_{k}]\preccurlyeq{\bm{\Upsilon}}_{2}$ almost surely, then the above inequality also holds, while we replace $\bm{M}$} by $\bm{W}$}, $\bm{X}$} by $\bm{D}$}, $d$ by $2[r({\bm{\Upsilon}}_{1}\wedge\|\|\left\langle{{\bm{W}}}\right\rangle_{\infty}\|\|_{\frac{p}{2}})\vee r({\bm{\Upsilon}}_{2}\wedge\|\|\left\langle{{\bm{W}}}\right\rangle_{\infty}\|\|_{\frac{p}{2}})]$, and double the right-hand side.
For self-adjoint matrix martingales $({\bm{M}}_{k})_{k\in{\mathbb{N}}}$, the proof strategy is to combine the extrapolation method (good $\lambda$ inequality) (Hitczenko, 1990, 1994; Pinelis, 1994) to derive a scalar-version Rosenthal-Burkholder inequality, with techniques in matrix concentration inequalities (Tropp, 2011, 2012, 2015). A detailed outline of the proof can be found in Section 4.1.
@@
It is worth noting that the assumption in Theorem (b), the quadratic variation $\langle{\bm{M}}\rangle_{\infty}$ has an almost sure upper bound $\bm{\Upsilon}$}, seems quite restrictive. However, such an assumption is exactly satisfied in deriving the dimension-free matrix Rosenthal inequality for uniformly geometrically ergodic Markov chains. We also note that doubling the r.h.s. in Theorem 2.1 (b) is not essential; we could reduce this multiplier from $2$ to a constant close to $1$ by increasing the effective rank term that only appears as the logarithmic term. However, for brevity, we use this presentation and do not strive for the optimal constants.
@@ Discussion.
In the i.i.d. case, Chen et al. (2012); Mackey et al. (2014) provided a matrix Rosenthal inequality, and Jirak et al. (2024) further gave a dimension-free version. We extend their results from the i.i.d. setting to the martingale setting. See also Junge and Xu (2003, 2008); Junge and Zeng (2013); Jiao et al. (2025) for the Rosenthal-Burkholder inequality and other probability inequalities for noncommutative martingales.
We also present another version of the Rosenthal-Burkholder inequality: $\|\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{k}\|\|_{p}\lesssim[(p/\log p)\vee\log d](\|\left\|{\left\langle{{\bm{M}}}\right\rangle_{\infty}}\right\|^{1/2}\|_{p}+\|\sup_{k\in{\mathbb{N}}}\|{\bm{X}}_{k}\|\|_{p})$, which can be found in Section 4.1.
@@ 2.2 Matrix Rosenthal and Concentration Inequalities for Markov Chains
Consider the Markov kernel ${\mathrm{Q}}\colon{\mathcal{Z}}\times{\mathscr{Z}}\to[0,1]$ on the measurable space $({\mathcal{Z}},{\mathscr{Z}})$ with a unique stationary distribution $\pi\in{\mathcal{P}}$, where $\mathcal{P}$} is the set of all probability distributions on $({\mathcal{Z}},{\mathscr{Z}})$. For any $\xi\in{\mathcal{P}}$, we denote by ${\mathbb{P}}_{\xi}$ (resp. ${\mathbb{E}}_{\xi}$) the probability (resp. expectation) of the canonical process $(Z_{k})_{k\in{\mathbb{N}}}$ with initial distribution $\xi$, and $({\mathscr{F}}_{k})_{k\in{\mathbb{N}}}$ the nature filtration where ${\mathscr{F}}_{k}=\sigma(Z_{0},\ldots,Z_{k})$. That is, $(Z_{k})_{k\in{\mathbb{N}}}$ is a stochastic process in $({\mathcal{Z}}^{\mathbb{N}},{\mathscr{Z}}^{\mathbb{N}},{\mathbb{P}}_{\xi})$ adapted to $({\mathscr{F}}_{k})_{k\in{\mathbb{N}}}$, which satisfies ${\mathbb{P}}_{\xi}(Z_{0}\in A)=\xi(A)$, ${\mathbb{P}}_{\xi}(Z_{k+1}\in A\mid{\mathscr{F}}_{k})={\mathbb{P}}_{\xi}(Z_{k+1}\in A\mid Z_{k})={\mathrm{Q}}(Z_{k},A)$ ${\mathbb{P}}_{\xi}$-almost surely for any $A\in{\mathscr{Z}}$ and $k\in{\mathbb{N}}$. Thus, $(Z_{k})_{k\in{\mathbb{N}}}$ is an Markov chain.
@@
Now, we introduce the ergodicity assumptions we consider in this paper, we will only consider the uniformly geometrically ergodic Markov chains in this section, and the more general geometrically $V$-ergodic Markov chains in Section 2.2.1. Under other ergodicity condition, such as geometrically ergodic Markov chains with respect to the Wasserstein semimetric (Durmus et al., 2023, Section 2.3), similar results can be derived by using the same techniques, so we omit them in this paper.
@@
$\sup_{z\in{\mathcal{Z}}}\left\|{{\mathrm{Q}}^{k}(z,\cdot)-\pi}\right\|_{\textup{TV}}\leq 2\left({1/4}\right)^{\lfloor k/t_{\operatorname{mix}}\rfloor},$
@@
We consider a measurable self-adjoint matrix function ${\bm{F}}\colon{\mathcal{Z}}\to{\mathbb{H}}^{d}$. We assume $\pi({\bm{F}})=\int_{{\mathcal{Z}}}{\bm{F}}(z)\pi(dz)=\bm{0}_{d{\times}d}$ and $\|\|{\bm{F}}\|\|_{\infty}=\sup_{z\in{\mathcal{Z}}}\|{\bm{F}}(z)\|<\infty$. Our goal is to give counterparts to matrix Rosenthal, Hoeffding, Bernstein inequalities for the spectral norm of the summation ${\bm{S}}_{n}=\sum_{k=0}^{n-1}{{\bm{F}}}(Z_{k})$. We observe that the boundedness condition is indispensable. As argued by Fan et al. (2021, Theorem 4), even for scalar unbounded functions, it is impossible to obtain Hoeffding inequalities for ${\bm{S}}_{n}$ under Assumption 2.1, whereas our Rosenthal inequality implies the Hoeffding-typ inequality for ${\bm{S}}_{n}$.
Recall that the Markov chain CLT (Jones, 2004, Corollary 5) says that $n^{-1/2}{\bm{S}}_{n}$ converges weakly to a Gaussian distribution and satisfies $\lim_{n\to\infty}n^{-1}{\mathbb{E}}_{\xi}[{\bm{S}}_{n}^{2}]={\bm{\Sigma}}_{\pi}({\bm{F}})$ for any $\xi\in{\mathcal{P}}$, where ${\bm{\Sigma}}_{\pi}({\bm{F}})$ is the long-run variance defined as
${\bm{\Sigma}}_{\pi}({\bm{F}}):=\pi\left({{{\bm{F}}}^{2}}\right)+\sum_{k=1}^{\infty}\pi\left({{{\bm{F}}}{\mathrm{Q}}^{k}{{\bm{F}}}+\left({{\mathrm{Q}}^{k}{{\bm{F}}}}\right){{\bm{F}}}}\right).$
We aim to obtain a Rosenthal inequality whose leading term is given by the long-run variance ${\bm{\Sigma}}_{\pi}({\bm{F}})$, matching the Markov chain CLT.
@@
Rectangular Case: In this case, we consider a rectangular matrix function ${\bm{R}}\colon{\mathcal{Z}}\to{\mathbb{C}}^{d_{1}{\times}d_{2}}$ that satisfies $\pi({\bm{R}})=\int_{{\mathcal{Z}}}{\bm{R}}(z)\pi(dz)=\bm{0}_{d_{1}{\times}d_{2}}$. We denote by ${\bm{V}}_{n}=\sum_{k=0}^{n-1}{\bm{R}}(Z_{k})$ the summation and
$\left\|{{\bm{\Sigma}}_{\pi}({\bm{R}})}\right\|:=\left\|{\pi\left({{\bm{R}}{\bm{R}}^{*}}\right){+}\sum_{k=1}^{\infty}\pi\left({{{\bm{R}}}{\mathrm{Q}}^{k}{{\bm{R}}^{*}}{+}\left({{\mathrm{Q}}^{k}{{\bm{R}}}}\right){{\bm{R}}^{*}}}\right)}\right\|\vee\left\|{\pi\left({{\bm{R}}^{*}{\bm{R}}}\right){+}\sum_{k=1}^{\infty}\pi\left({{{\bm{R}}^{*}}{\mathrm{Q}}^{k}{{\bm{R}}}{+}\left({{\mathrm{Q}}^{k}{{\bm{R}}^{*}}}\right){{\bm{R}}}}\right)}\right\|,$ (7)
@@
$\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\pi,p}\leq$ $C_{R,1}\sqrt{\frac{(p\vee\log d)\left\|{{\bm{\Sigma}}_{\pi}({\bm{F}})}\right\|}{n}}+D_{1}t_{\operatorname{mix}}^{1/2}\left({\frac{(p\vee\log d)t_{\operatorname{mix}}}{n}}\right)^{3/4}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty},$
$\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\xi,p}\leq\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\pi,p}+D_{2}\frac{pt_{\operatorname{mix}}}{n}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty},$ (8)
where $C_{R,1}$ is defined in Theorem 2.1 and $D_{1},D_{2}>0$ are universal constants, which can be traced in the proof.
@@
    Dimension-Free: in the self-adjoint case, if we further assume that there exists a ${\bm{\Upsilon}}\in{\mathbb{H}}^{d}_{+}$ such that ${\bm{F}}^{2}(z)\preccurlyeq{\bm{\Upsilon}}$, then the above inequality also holds when we replace the $d$ in the first term by $e\cdot r([(64/9)t_{\operatorname{mix}}^{2}{\bm{\Upsilon}}]\wedge[(1/n)\|\|\left\langle{{\bm{M}}}\right\rangle_{n}\|\|_{\pi,{p}/{2}}])$, and replace the $d$ in the second term by the dimension factor defined after Eqn. (17), and double the right-hand side.

### v2

@@ 2 Main Results
In this section, we present our main results: the Rosenthal-Burkholder inequality for discrete-time matrix martingales and Burkholder-Davis-Gundy inequality for continuous matrix local martingales in Section 2.1, as well as matrix Rosenthal, Hoeffding, and Bernstein inequalities for ergodic Markov chains in Section 2.2. We will also give dimension-free versions of these results.
We find that the dimension-free results can be applied to linear operator-valued martingales in infinite-dimensional separable Hilbert spaces. For brevity, we will not change the notation, and our dimension-free results still hold if ${\mathbb{H}}^{d}$ is replaced by the space of all self-adjoint operators in the separable Hilbert space $\mathcal{X}$}, and ${\mathbb{C}}^{d_{1}{\times}d_{2}}$ is replaced by the space of all bounded linear operators from the separable Hilbert space ${\mathcal{X}}_{1}$ to the separable Hilbert space ${\mathcal{X}}_{2}$.
@@ 2.1 The Rosenthal-Burkholder Inequality for Matrix Martingales
We denote by $({\bm{X}}_{k})_{k\in{\mathbb{N}}}$ the martingale difference sequence, that is, ${\bm{X}}_{k}={\bm{M}}_{k}-{\bm{M}}_{k-1}\in{\mathbb{H}}^{d}$ (${\bm{M}}_{-1}:=\bm{0}_{d{\times}d}$). And we denote by $(\left\langle{{\bm{M}}}\right\rangle_{k})_{k\in{\mathbb{N}}}$ the predictable quadratic variation sequence, that is, $\left\langle{{\bm{M}}}\right\rangle_{k}=\sum_{j=0}^{k}{\mathbb{E}}_{j-1}{\bm{X}}_{j}^{2}\in{\mathbb{H}}^{d}_{+}$. Throughout this section, we assume that $\left\langle{{\bm{M}}}\right\rangle_{\infty}:=\lim_{k\to\infty}\left\langle{{\bm{M}}}\right\rangle_{k}<\infty$ element-wisely almost surely. As a result, by the martingale convergence theorem (Durrett, 2019, Theorem 4.5.2), $\lim_{k\to\infty}{{\bm{M}}}_{k}={\bm{M}}_{\infty}\in{\mathbb{H}}^{d}$ almost surely.
@@
Dimension-Free: As mentioned in Minsker (2017), in the case of ultra-high dimensions or even infinite dimensions, it is desirable that the results do not depend on the dimension $d$ of the ambient space. To this end, we resort to the notion of effective rank (Koltchinskii and Lounici, 2016). For any ${\bm{\Upsilon}}\in{\mathbb{H}}^{d}_{+}$, its effective rank is defined as
[…]
It is easily seen that $r({\bm{\Upsilon}})\leq d$ and $r({\bm{\Upsilon}})$ can be much smaller than $d$ if the eigenvalues of $\bm{\Upsilon}$} rapidly decay to 0.
@@
$\left\|{\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{k}\|}\right\|_{p}\leq C_{1}\sqrt{p\vee\log d}\Big\|\left\|{\left\langle{{\bm{M}}}\right\rangle_{\infty}}\right\|^{1/2}\Big\|_{p}+C_{2}(p\vee\log d)\left\|{\sup_{k\in{\mathbb{N}}}\|{\bm{X}}_{k}\|}\right\|_{p},$ (8)
where $C_{1}=87$, $C_{2}=50$. When $p\geq 117$, the constants can be smaller: $C_{1}=64$, $C_{2}=28$.
@@
    In the rectangular case (a), if we further assume that there exist ${\bm{\Upsilon}}_{1}\in{\mathbb{H}}_{+}^{d_{1}}$, ${\bm{\Upsilon}}_{2}\in{\mathbb{H}}_{+}^{d_{2}}$ such that $\sum_{k=0}^{\infty}{\mathbb{E}}_{k-1}[{\bm{D}}_{k}{\bm{D}}_{k}^{*}]\preccurlyeq{\bm{\Upsilon}}_{1}$ and $\sum_{k=0}^{\infty}{\mathbb{E}}_{k-1}[{\bm{D}}_{k}^{*}{\bm{D}}_{k}]\preccurlyeq{\bm{\Upsilon}}_{2}$ almost surely, then the above inequality also holds, while we replace $\bm{M}$} by $\bm{W}$}, $\bm{X}$} by $\bm{D}$}, $d$ by $2[r({\bm{\Upsilon}}_{1}\wedge\|\|\left\langle{{\bm{W}}}\right\rangle_{\infty}\|\|_{\frac{p}{2}})\vee r({\bm{\Upsilon}}_{2}\wedge\|\|\left\langle{{\bm{W}}}\right\rangle_{\infty}\|\|_{\frac{p}{2}})]$, and double the right-hand side.
By localization, our Rosenthal-Burkholder inequalities also hold if we consider matrix local martingales and replace $k\in{\mathbb{N}}$ by $0\leq k\leq\tau$ for any stopping time $\tau$; we omit the details for brevity. We also present another version of the Rosenthal-Burkholder inequality: $\|\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{k}\|\|_{p}\lesssim[(p/\log p)\vee\log d](\|\left\|{\left\langle{{\bm{M}}}\right\rangle_{\infty}}\right\|^{1/2}\|_{p}+\|\sup_{k\in{\mathbb{N}}}\|{\bm{X}}_{k}\|\|_{p})$, which can be found in Section 4.1.
For self-adjoint matrix martingales $({\bm{M}}_{k})_{k\in{\mathbb{N}}}$, the proof strategy is to combine the extrapolation method (good $\lambda$ inequality) (Hitczenko, 1990, 1994; Pinelis, 1994) with techniques in matrix concentration inequalities (especially matrix Bennett inequalities) (Tropp, 2011, 2012, 2015; Kroshnin and Suvorikova, 2024). A detailed outline of the proof can be found in Section 4.1.
@@
It is worth noting that the assumption in Theorem 2.1 (b), the quadratic variation $\langle{\bm{M}}\rangle_{\infty}$ has an almost sure upper bound $\bm{\Upsilon}$}, seems quite restrictive. However, such an assumption is exactly satisfied in deriving the dimension-free matrix Rosenthal inequality for uniformly geometrically ergodic Markov chains.
We also note that doubling the r.h.s. in Theorem 2.1 (b) is not essential; we could reduce this multiplier from $2$ to a constant close to $1$ by increasing the effective rank term that only appears as the logarithmic term. However, for brevity, we use this presentation and do not strive for the optimal constants.
A direct application of the Rosenthal-Burkholder inequality (Theorem 2.1) is to show the Burkholder-Davis-Gundy inequality for continuous matrix local martingales by discretization in the following theorem.
Theorem 2.2 (Burkholder-Davis-Gundy Inequality).
Let $({\bm{M}}_{t})_{t\geq 0}$ be a continuous local martingale taking values in ${\mathbb{H}}^{d}$. We denote by $(\left\langle{{\bm{M}}}\right\rangle_{t})_{t\geq 0}$ its quadratic variation process, that is, $\left\langle{{\bm{M}}}\right\rangle_{t}]_{ij}=\sum_{k=1}^{d}\left\langle{[{\bm{M}}]_{ik},[{\bm{M}}]_{kj}}\right\rangle_{t}$. For any $p\geq 2$ and stopping time $\tau$, suppose that $\|\left\|{\left\langle{{\bm{M}}}\right\rangle_{\tau}}\right\|^{1/2}\|_{p}<\infty$. Then it holds that
$\left\|{\sup_{0\leq t\leq\tau}\|{\bm{M}}_{t}\|}\right\|_{p}\leq C_{1}\sqrt{p\vee\log d}\Big\|\left\|{\left\langle{{\bm{M}}}\right\rangle_{\tau}}\right\|^{1/2}\Big\|_{p},$
where $C_{1}$ is defined in Theorem 2.1.
The proof can be found in Appendix A.2. Maitre (2025, Theorem 1.1) also provided a version of Burkholder-Davis-Gundy inequality for continuous matrix martingales in the form: $\|\sup_{0\leq t\leq\tau}\|{\bm{M}}_{t}\|\|_{p}\lesssim\sqrt{p}\|\left\|{\left\langle{{\bm{M}}}\right\rangle_{\tau}}\right\|\|_{p}^{1/2}+\sqrt{\log d}\|\left\|{\left\langle{{\bm{M}}}\right\rangle_{\tau}}\right\|^{1/2}\|_{p}$, our version (Theorem 2.2) is sharper because $\|\left\|{\left\langle{{\bm{M}}}\right\rangle_{\tau}}\right\|^{1/2}\|_{p}\leq\|\left\|{\left\langle{{\bm{M}}}\right\rangle_{\tau}}\right\|\|_{p}^{1/2}$ by Jensen’s inequality.
@@ Discussion.
In the independent case, Chen et al. (2012, Theorem A.1) provided a matrix Rosenthal inequality, and Jirak et al. (2025, Theorem 3) further gave a dimension-free version. We extend their results from the independent case to the martingale case. We remark that, in the independent case, it is possible to translate the universality-based concentration inequalities (Brailovskaya and van Handel, 2024, Theorem 2.8 and Theorem 2.16) into Rosenthal-type inequalities by integrating the tail bounds. In some cases, this yields tighter upper bounds than the classic results Chen et al. (2012, Theorem A.1). This observation gives rise to an interesting question: in the martingale case, does certain form of universality-based moment inequality hold? We leave this as a direction for future research.
In the field of noncommutative probability, Junge and Xu (2003, 2008); Junge and Zeng (2013, 2015); Jiao et al. (2025) investigated probability inequalities for noncommutative martingales. In our random matrix setting, their results can be translated into probability inequalities concerning the Schatten $p$-norm, rather than the spectral norm that is the focus of this paper. In particular, in the independent case, Junge and Zeng (2013, Theorem 0.4) gives the following Rosenthal inequality
${\mathbb{E}}^{{1}/{p}}\left[{\operatorname{Tr}\left({\left|{{\bm{M}}_{\infty}}\right|^{p}}\right)}\right]\lesssim\sqrt{p}{\mathbb{E}}^{{1}/{p}}\left[{\operatorname{Tr}\left({\left\langle{{\bm{M}}}\right\rangle_{\infty}^{p/2}}\right)}\right]+p{\mathbb{E}}^{{1}/{p}}\left[{\sup_{k\in{\mathbb{N}}}\operatorname{Tr}\left({\left|{{\bm{X}}_{k}}\right|^{p}}\right)}\right],$
for $p\geq 2.5$. However, in the martingale case, whether a Rosenthal-Burkholder inequality of the following form holds remains an open problem (Junge and Zeng, 2015, Problem 2.5)
${\mathbb{E}}^{{1}/{p}}\left[{\operatorname{Tr}\left({\left|{{\bm{M}}_{\infty}}\right|^{p}}\right)}\right]\lesssim\sqrt{p}{\mathbb{E}}^{{1}/{p}}\left[{\operatorname{Tr}\left({\left\langle{{\bm{M}}}\right\rangle_{\infty}^{p/2}}\right)}\right]+f(p){\mathbb{E}}^{{1}/{p}}\left[{\sum_{k\in{\mathbb{N}}}\operatorname{Tr}\left({\left|{{\bm{X}}_{k}}\right|^{p}}\right)}\right],$
where the coefficient of the dominant term (quadratic variation) has an optimal $\sqrt{p}$ dependence on $p$, and $f(p)$ is some function for which they conjectured $f(p)=p$ as in the commutative case.
@@ 2.2 Matrix Rosenthal and Concentration Inequalities for Markov Chains
Consider the Markov kernel ${\mathrm{Q}}\colon{\mathcal{Z}}\times{\mathscr{Z}}\to[0,1]$ on the measurable space $({\mathcal{Z}},{\mathscr{Z}})$ with a unique stationary distribution $\mu\in{\mathcal{P}}$, where $\mathcal{P}$} is the set of all probability distributions on $({\mathcal{Z}},{\mathscr{Z}})$. For any $\xi\in{\mathcal{P}}$, we denote by ${\mathbb{P}}_{\xi}$ (resp. ${\mathbb{E}}_{\xi}$) the probability (resp. expectation) of the canonical process $(Z_{k})_{k\in{\mathbb{N}}}$ with initial distribution $\xi$, and $({\mathscr{F}}_{k})_{k\in{\mathbb{N}}}$ the nature filtration where ${\mathscr{F}}_{k}=\sigma(Z_{0},\ldots,Z_{k})$. That is, $(Z_{k})_{k\in{\mathbb{N}}}$ is a stochastic process in $({\mathcal{Z}}^{\mathbb{N}},{\mathscr{Z}}^{\mathbb{N}},{\mathbb{P}}_{\xi})$ adapted to $({\mathscr{F}}_{k})_{k\in{\mathbb{N}}}$, which satisfies ${\mathbb{P}}_{\xi}(Z_{0}\in A)=\xi(A)$, ${\mathbb{P}}_{\xi}(Z_{k+1}\in A\mid{\mathscr{F}}_{k})={\mathbb{P}}_{\xi}(Z_{k+1}\in A\mid Z_{k})={\mathrm{Q}}(Z_{k},A)$ ${\mathbb{P}}_{\xi}$-almost surely for any $A\in{\mathscr{Z}}$ and $k\in{\mathbb{N}}$. Thus, $(Z_{k})_{k\in{\mathbb{N}}}$ is an Markov chain.
@@
Now, we introduce the ergodicity assumptions we consider in this paper, we will only consider the uniformly geometrically ergodic Markov chains in this section, and the more general geometrically $V$-ergodic Markov chains in Section 2.2.1. We recall that an irreducible and aperiodic Markov chain with a finite state space is always uniformly geometrically ergodic.
@@
$\sup_{z\in{\mathcal{Z}}}\left\|{{\mathrm{Q}}^{k}(z,\cdot)-\mu}\right\|_{\textup{TV}}\leq 2\left({1/4}\right)^{\lfloor k/t_{\operatorname{mix}}\rfloor},$
@@
We consider a measurable self-adjoint matrix function ${\bm{F}}\colon{\mathcal{Z}}\to{\mathbb{H}}^{d}$. We assume $\mu({\bm{F}})=\int_{{\mathcal{Z}}}{\bm{F}}(z)\mu(dz)=\bm{0}_{d{\times}d}$ and $\|\|{\bm{F}}\|\|_{\infty}=\sup_{z\in{\mathcal{Z}}}\|{\bm{F}}(z)\|<\infty$. Our goal is to give counterparts to matrix Rosenthal, Hoeffding, Bernstein inequalities for the spectral norm of the summation ${\bm{S}}_{n}=\sum_{k=0}^{n-1}{{\bm{F}}}(Z_{k})$. We observe that the boundedness condition is indispensable. As argued by Fan et al. (2021, Theorem 4), even for scalar unbounded functions, it is impossible to obtain Hoeffding-type inequalities for ${\bm{S}}_{n}$ under Assumption 2.1 or a non-zero absolute $L^{2}$-spectral gap, whereas our Rosenthal inequality implies the Hoeffding-type inequality for ${\bm{S}}_{n}$.
Recall that the Markov chain CLT (Jones, 2004, Corollary 5) says that $n^{-1/2}{\bm{S}}_{n}$ converges weakly to a Gaussian distribution and satisfies $\lim_{n\to\infty}n^{-1}{\mathbb{E}}_{\xi}[{\bm{S}}_{n}^{2}]={\bm{\Sigma}}_{\mu}({\bm{F}})$ for any $\xi\in{\mathcal{P}}$, where ${\bm{\Sigma}}_{\mu}({\bm{F}})$ is the long-run variance defined as
${\bm{\Sigma}}_{\mu}({\bm{F}}):=\mu\left({{{\bm{F}}}^{2}}\right)+\sum_{k=1}^{\infty}\mu\left({{{\bm{F}}}{\mathrm{Q}}^{k}{{\bm{F}}}+\left({{\mathrm{Q}}^{k}{{\bm{F}}}}\right){{\bm{F}}}}\right).$
We aim to obtain a Rosenthal inequality whose leading term is given by the long-run variance ${\bm{\Sigma}}_{\mu}({\bm{F}})$, matching the Markov chain CLT.
@@
Rectangular Case: In this case, we consider a rectangular matrix function ${\bm{R}}\colon{\mathcal{Z}}\to{\mathbb{C}}^{d_{1}{\times}d_{2}}$ that satisfies $\mu({\bm{R}})=\int_{{\mathcal{Z}}}{\bm{R}}(z)\mu(dz)=\bm{0}_{d_{1}{\times}d_{2}}$. We denote by ${\bm{V}}_{n}=\sum_{k=0}^{n-1}{\bm{R}}(Z_{k})$ the summation and
[…]
@@
$\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\mu,p}\leq$ $C_{1}\sqrt{\frac{(p\vee\log d)\left\|{{\bm{\Sigma}}_{\mu}({\bm{F}})}\right\|}{n}}+D_{1}t_{\operatorname{mix}}^{1/2}\left({\frac{(p\vee\log d)t_{\operatorname{mix}}}{n}}\right)^{3/4}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty},$
$\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\xi,p}\leq\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\mu,p}+D_{2}\frac{pt_{\operatorname{mix}}}{n}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty},$ (10)
where $C_{1}$ is defined in Theorem 2.1 and $D_{1},D_{2}>0$ are universal constants, which can be traced in the proof.
@@
    Dimension-Free: in the self-adjoint case, if we further assume that there exists a ${\bm{\Upsilon}}\in{\mathbb{H}}^{d}_{+}$ such that ${\bm{F}}^{2}(z)\preccurlyeq{\bm{\Upsilon}}$, then the above inequality also holds when we replace the $d$ in the first term by $d_{n,p}=e\cdot r([(64/9)t_{\operatorname{mix}}^{2}{\bm{\Upsilon}}]\wedge[(1/n)\|\|\left\langle{{\bm{M}}}\right\rangle_{n}\|\|_{\mu,{p}/{2}}])$, and replace the $d$ in the second term by the dimension factor defined after Eqn. (19), and double the right-hand side.